OpenStudy (anonymous):

Suppose L: R^3 maps R^3 is a matrix transformation. and L((0,2,-1))=(-1,1,-4) and L((0,1,-1))=(1,-5,5). Find L(3(0,2,-1)-2(0,1,-1) = L((0,-5,3)) = Help please!

OpenStudy (amistre64):

L is linear, so what does that tell us about scalars?

OpenStudy (anonymous):

There scalable?

OpenStudy (amistre64):

L(kx) for some constant k is equal to: kL(x)

OpenStudy (anonymous):

Right, that's the formula for matrix transformations. Not sure how to proceed with it to solve the problem

OpenStudy (amistre64):

also, a linear property is such that L(x+y) = L(x)+L(y), putting this to your problem we get: L(3(0,2,-1)-2(0,1,-1) L(3(0,2,-1)) + L(-2(0,1,-1) ) 3 L(0,2,-1) -2 L(0,1,-1)

OpenStudy (amistre64):

and they define the L parts for those ...

OpenStudy (amistre64):

now, a trickier part may be to find 0,-5,3 as a linear combination of the vectors being Ld

OpenStudy (anonymous):

How did you get from the 2nd one to the last one?

OpenStudy (amistre64):

pulled out the scalars of course

OpenStudy (anonymous):

Oh yeah i see it now, but that's all there is to the solution? I thought it would be a lot harder

OpenStudy (amistre64):

you have presented to you: L(x) = u, L(y) = v, and for some sclaras, j,k L(kx + jy) = ?? L(kx) + L(jy) kL(x) + jL(y) ku + jv

OpenStudy (amistre64):

now the question becomes trickier for the 2nd one: find kx + jy = (0,-5,3) and then apply the operations from the first part

OpenStudy (anonymous):

Awesome it became a lot easier to understand, thanks!

OpenStudy (amistre64):

youre welcome :)